OpenStudy (anonymous):

k^3(k ^7/5)^-5 What is the simplified form of the expression?

OpenStudy (mertsj):

\[k^3(\frac{k^7}{5})^{-5}=k^3(\frac{5}{k^7})^{5}=k^3(\frac{3125}{k ^{35}})=\frac{3125}{k ^{32}}\]